sarah is 15 yrs older than joseph. in ten years from now. sarah will be twice as old as joseph.
how old is sarah now?

Answer by ankor@dixie-net.com(22740)

sarah is 15 yrs older than joseph.
s = j+15
in ten years from now. sarah will be twice as old as joseph.
s+10 = 2(j+10)
s + 10 = 2j + 20
s = 2j + 20 - 10
s = 2j + 10
Replace s with (j+15)
j + 15 = 2j + 10
15 - 10 2j - j
j = 5 yrs for Joseph
:
:
how old is sarah now? I'll let you figure that out
:
Check the solutions in the 2nd statement